wolfheart:

The map of a walking trail is drawn on a coordinate grid with three points of interest. The trail starts at R(−1, 4) and goes to S(5, 4) and continues to T(5, −2). The total length of the walking trail is ____ units. (Input whole numbers only.)

supie:

\(\bf\ First\ we\ have\ to\ know\ the\ the\ formula\ for\ distance \) \(\sf\ Distance\ Formula:\) \[D=\sqrt{\left(x_2-x_1\right)^2+\left(y_2-y_1\right)^2}\]

supie:

\[\bf\ RS + ST\] is to get the `total trail` So we have to find the ordered pairs of the variables, R, S, and T \(R(-1,\ 4)\ S(5,\ 4)\ T(5,\ -2\))

supie:

So, we have \[RS=\sqrt{\left(5-\left(-1\right)\right)^2+\left(4-4\right)^2}\]

supie:

So basically, the fastest and easy way would be \[RS=\sqrt{\left(5-\left(-1\right)\right)^2+\left(4-4\right)^2}\]\[RS=√(5+1)^2\]\[RS=√6^2\]\[RS=√(2*3)^2\]RS=2^2/2*3^2/2 = RS=2*3 = RS=6 \(\frac{RS}{S}=\frac{6}{s}\)=\(R=\frac{6}{s}\)\(\sf\ meaning\ that\ RS=6\)

supie:

Then do the same for the second one and add the two end trails together then that will give you your answer :)